# Prompting Large Language Models for Recommender Systems: A Comprehensive Framework and Empirical Analysis

### v1

@@ 5.1. Task Description
Task description of recommendation tasks for ProLLM4Rec is a key part of prompts that makes LLM understand the goal of specific recommendation assignments. Typically, based on the different recommendation objectives, existing recommendation tasks can be divided into the following categories: point-wise optimization for a single item (e.g., binary cross entropy loss (Rendle, 2010)), pair-wise optimization for positive and negative items (e.g., Bayesian personalized ranking loss (Rendle et al., 2009)), and list-wise optimization for multiple items (e.g., variational auto-encoders (Liang et al., 2018)). For ProLLM4Rec, we can adapt our work into multiple tasks by changing the description of prompts and the form of candidate items. That is to say, our framework can not only be used for re-ranking (Hou et al., 2023) and retrieving (Li et al., 2023b) tasks, but also support task forms such as point-wise prediction (Kang et al., 2023), pair-wise comparison (Dai et al., 2023a) and list-wise re-ranking formats.
$\bullet$ Point-wise recommendation  (Cheng et al., 2016) regards the recommendation task as a binary classification problem, such as the rating scoring task (Geng et al., 2022; Chu et al., 2023) and click-through rate (CTR) predictions (Kang et al., 2023; Bao et al., 2023b). It generates a score or likelihood of preference for a given user-item pair with feature engineering. The recommender system ranks items solely based on the individual characteristics or historical interactions of users without considering the comparison and mutual influence between items in the input ranking list. When utilizing LLMs for point-wise recommendation, the description mainly involves the user-item information with the range of score or answer list to limit output formats of LLMs.
@@
$\bullet$ List-wise recommendation (Liang et al., 2018) involves optimizing an entire list of recommended items for a user. Instead of considering items individually or in pairs, this method treats the entire recommendation list as a single entity and captures the interior correlations within the list. It aims to create a ranked list of items that collectively maximizes user satisfaction. For LLM-based list-wise recommendations, the prompt contains a list of items and corresponding instructions, eliciting the ability of LLMs to explore the potential relationship within the item sequence (Hou et al., 2023; Ma et al., 2023).
$\bullet$ Matching (Bao et al., 2023a; Li et al., 2023b) in recommendation involves obtaining a small subset of candidates from the entire item pool for personalized users, and candidate items are not provided in prompts at this stage. Due to the lack of domain knowledge in recommender systems, leveraging LLMs for item retrieval requires additional strategies such as model training, item indexing and grounding approaches (Zheng et al., 2023). The prompt for matching tasks includes the description of a user in different forms, e.g., textual profiles and user-item interaction histories, in order to quickly match suitable items.
$\bullet$ Ranking  (Luo et al., 2023b) in recommendation involves sorting or ordering items based on the predicted relevance or probability of interest to the target user. It aims to present the most relevant items at the top of the recommendation list. By analyzing historical user-item interactions, user preferences, item features, and the given cadidate item list, it requires LLMs to estimate the relevance of items and determine corresponding positions in the recommendation list (Li et al., 2023k; Luo et al., 2023b; Gao et al., 2023).
@@ 5.2. User Interest Modeling
Compared to general tasks solved by LLMs, the characteristic of recommendation tasks lies in the mining and utilization of the user personalized interest (Yao et al., 2023; Shu et al., 2023). Users are influenced by multiple factors when making recommendation decisions, including but not limited to their long-term preferences, short-term intentions, market popular tendencies, occasional environmental biases, and so on (Cheng et al., 2016; He et al., 2017). Among them, the most essential aspect is the long-term preferences and short-term intentions of users, which is the purpose of recommender systems (He et al., 2020; Zhou et al., 2020; Hou et al., 2022). In this section, we first classify the user interest type (Section 5.2.1), outline the user representation forms (Section 5.2.2), and summarize existing modeling methods (Section 5.2.3). Then, we discuss approaches to modeling short-term (Section 5.2.4) and long-term interest (Section 5.2.5) of users, and conduct experiments to provide empirical analysis and key findings.
@@ 5.2.1. User Interest Type
In the field of recommender systems, there are multiple ways to classify the user interest (Rendle et al., 2009; Guo et al., 2017; Zhou et al., 2020). For example, based on the type of feedback between users and the platform, the user interest can be divided into the explicit interest (e.g., ratings and reviews) and implicit interest (e.g., clicks). Although the explicit feedback can reflect the true intentions of users, the sparse and expensive data limits its application scenarios. In general, we mainly explore user interest modeling in LLMs under implicit feedback scenarios. Another classification of the user interest is based on the time duration and stability, i.e., short-term intentions, long-term preferences, and hybrid interest as follows:
$\bullet$ Short-term intentions refer to the sudden and accidental intentions or tendencies of personalized users in recent interactions, which are prone to change and can be influenced by environmental factors in recommender systems (Lin et al., 2023b; Hou et al., 2023). That is to say, short-term intentions are recent, temporary, and variable (Zhang et al., 2023f).
$\bullet$ Long-term preferences refer to the stable preferences of users towards certain content, themes, and elements, which are not easily changed in the short term and will continue to affect recommendation decisions of users (Shu et al., 2023; Wang et al., 2023c). In contrast to short-term intentions, long-term preferences are long-term, sustained and stable.
$\bullet$ Hybrid interest refers to the combination of long-term preferences and short-term intentions (Friedman et al., 2023; Liu et al., 2023e; Zhiyuli et al., 2023). On the one hand, short-term intentions are dominated by long-term preferences (Yue et al., 2023). On the other hand, long-term preferences also consist of the short-term interest across multiple time periods (Li et al., 2023j; Qiu et al., 2023).
@@ 5.2.2. User Representation Forms
In ProLLM4Rec, user interest modeling should consider both the input form for the prompt template and the specific storage form for the interest memory (Wang et al., 2023c). Generally, there are three types of input contents for interest modeling in LLMs: historical item lists, interest descriptions and user embeddings. Corresponding to different representation forms of the user interest, there are also various storage forms in the interest memory.
@@
$\bullet$ Historical item lists refer to representing personalized users based on the sequence of historical interacted items, which is widely used in session-based recommendation and sequential recommendation (Hou et al., 2022; Zhou et al., 2020; Kang and McAuley, 2018). For token-based item indexing, the ID sequence of items is used for user modeling in LLMs, similar to traditional sequential recommendation models (Geng et al., 2022). However, without fine-tuning, LLMs cannot recognize the meaning of corresponding IDs. For description-based item indexing, existing researches generally concatenate the attributes (e.g., titles) of items in the temporal order, and then input the item sequence as text descriptions to LLMs as user representations (Li et al., 2023d; Zhiyuli et al., 2023; Hou et al., 2023). Only the item IDs that the user has interacted with need to be stored as the interest memory. The list of items can be stored in the form of numpy or tensor arrays. Despite the simplicity and effectiveness, text attributes such as titles cannot fully represent items with ambiguity (Li et al., 2023d). At the same time, item sequences inevitably contain noise information (Wei et al., 2023), and limited sequence lengths make it difficult to accurately express the user interest based on pure lists.
@@
$\bullet$ Interest descriptions refer to representing users based on textual descriptions, which is more applicable to the text input of LLMs (Yao et al., 2023). That is to say, we can use natural languages to model the user interest in the form of text and input them into LLMs (Xi et al., 2023). To facilitate the storage and retrieval of textual descriptions, vector stores are commonly used (Touvron et al., 2023b). In LLMs, vector stores in the form of key-value pairs can connect implicit vector embeddings with explicit textual descriptions, improving the retrieval efficiency and comprehension ability (Li et al., 2023b; Petrov and Macdonald, 2023). However, the key lies in how to mine and describe the user interest, and this is what we should make efforts to research and discuss.
@@
$\bullet$ User embeddings refer to concatenating embeddings of users to the input of LLMs as user representations (Zhang et al., 2023b), which is often used for efficient fine-tuning of language models. In this case, we assign each user a unique ID and corresponding embedding, and add the vector representations of users to the input of LLMs for fine-tuning. At the same time, user embeddings can also be trained from small-scale traditional recommendation models (Zhang et al., 2023d), thereby incorporating collaborative filtering features from other users. However, it is worth noting that vector embeddings of user representations are only suitable for scenarios where the domain knowledge can be injected into LLMs, and the black-box property of embeddings also poses challenges to the explainability and understanding of language models.
@@ 5.2.3. Modeling Methods
In line with existing literature (Zhao et al., 2023; Wang et al., 2023g), we classify the methods for modeling interest in LLMs into three categories, i.e., memory-based methods, retrieval-based methods and generation-based methods.
@@
$\bullet$ Memory-based methods assign external memory to users for storing interest-related historical information (Xi et al., 2023; Shu et al., 2023). As for user representations, the encoded interest of users can be obtained from the memory (Zhang et al., 2023d), and LLMs are instructed to utilize the memorized interest with carefully designed prompts. It can be seen that the iterative updating of memories is the key to memory-based methods. As for personalized memory, there are three important operations (Wang et al., 2023g): (1) memory reading is to obtain contents of the interest memory based on the user identifier. (2) Memory writing is to augment new interest of users based on the latest interactions between users and items. (3) Memory reflection is the periodic examination and updating of existing issues in the memory. Strategies for memory reflection include but are not limited to self-summarization, self-correction, and reflection based on user feedback (Wang et al., 2023a). It is memory reflection that makes memory not just a stack of historical records but a summary of user interest (Yao et al., 2023). Note that memory-based methods specifically refer to obtaining contents from the memory without further processing procedures.
@@
$\bullet$ Retrieval-based methods add a module for personalized interest retrieval on top of the memory-based methods, utilizing the personalized query and retrieval strategies to obtain user interest from the personalized memory (Wang et al., 2023c; Lin et al., 2023b; Huang et al., 2023). As for the personalized query, there are generally three ways to form a query for retrieval: (1) candidate items as search queries (Lin et al., 2023b) for relevant items, (2) recently interacted items as search queries for similar items (Zhang et al., 2023c), and (3) the user profile as search queries for personalized items (Wang et al., 2023c). For the criteria of retrieval, there are generally multiple trade-offs, including dimensions such as relevance, recency, and diversity with respect to the user interest (Wang et al., 2023g).
@@
$\bullet$ Generation-based methods utilize the understanding and generation capabilities of LLMs to summarize, infer, and derive comprehensive user interest (Yao et al., 2023; Wang and Lim, 2023; Luo et al., 2023b). Generally, a combination of memory-based and retrieval-based methods is required for generation-based methods. For generative language models, the user interest induced by generative retrieval can further stimulate the prompting ability of LLMs (Liu et al., 2023a; Zhang et al., 2023e). The generative approach can also be used for data augmentation, leveraging the general knowledge of LLMs to augment textual descriptions of the user interest. In current recommendations, LLMs mainly utilize the retrieved relevant items to generate summarized interest descriptions, which can also serve as a strategy for the memory reflection (Wang et al., 2023c; Shu et al., 2023).
5.2.4. Research Questions and Observations for Modeling Short-term Intentions
@@
As discussed in Section 5.2.1, the user interest modeling in prompts for LLMs can be divided into: (1) short-term intentions, and (2) long-term preferences. In this section, we explore methods for modeling short-term intentions w.r.t. the following two research questions.
  • RQ1: How to select recent and relevant items for modeling short-term intentions?
  • RQ2: Whether different modeling forms of short-term intentions have similar recommendation performance?
@@
To analyze how to select recent and relevant items as the short-term intentions, we adopt the retrieval-based strategy to retrieve items from the long-term interest. Specifically, following the memory-based and retrieval-based strategy, the personalized query, e.g., candidate items, recently interacted items or the user profile, is encoded into vectors in a unified form. The most relevant items are retrieved from memory based on semantic similarity. Meanwhile, we set weights based on the recency of interactions so that the recent items have a higher probability of being retrieved. Finally, the corresponding retrieved text is obtained based on key-value pairs in the memory. Here, we explore the design of the personalized query and memory of users for item retrieval, and study the influence of different variants as follows.
@@
  • Variants of the personalized query: as for the personalized query, we utilize the interest of users for retrieving items, and design the following two variants to compare strategies for constructing the query.
@@
    • Long- and short-term interest: we first summarize the personalized profile of each user based on all items he or she has interacted, then concatenate the user profile and the summarized text based on 10 recently interacted items for the memory retrieval of the long-term interest.
  • Variants of the memory: as for the personalized memory, we consider two variants of the memory contents.
    • Global memory: we utilize descriptions of items in the datasets as the memory. Since all users in a dataset share the same item space, the memory w.r.t. item descriptions is not the personalized memory but global memory, indicating that different users store the same content for the same item.
@@
When the items utilized to recent interest modeling are fixed, different representation forms of the items may also affect the modeling results. To examine the impact of modeling forms on the short-term interest, we consider the recently interacted items without retrieval methods, and design the following four variants of short-term intentions:
  • Recent items: we use titles of the recent 10 items as short-term intentions. It is the default and baseline strategy.
  • Recent items + summarized text on recent items: we utilize the least-to-most prompting strategy as mentioned in Section 5, which summarizes the interest text based on the recently interacted 10 items, then concatenate the summarized text and the 10 items for modeling short-term intentions.
  • Personalized text of recent items: we construct the personalized memory for each user based on the user-item interactions, and use the personalized descriptions of recently interacted items as short-term intentions.
  • Recent items + personalized text of recent items: we employ not only titles of the recent 10 items, but also the personalized text of recent items to prompt LLMs as the short-term interest.
Table 6. Performance comparison on the design of personalized query and memory for item retrieval.
query memory MovieLens-1M Amazon-Books
ndcg@1 ndcg@10 ndcg@1 ndcg@10
@@
long- and short-term interest global memory 0.2200 0.4473 0.2400 0.3980
short-term interest personalized memory 0.2400 0.4441 0.3250 0.4636
@@
Table 7. Performance comparison on various modeling forms of short-term interest.
Short-term interest MovieLens-1M Amazon-Books
ndcg@1 ndcg@10 ndcg@1 ndcg@10
recent items 0.1817 0.3985 0.2467 0.4276
recent items + summarized text on recent items 0.2783 0.5099 0.2617 0.4729
personalized text of recent items 0.2100 0.4380 0.2500 0.4733
recent items + personalized text of recent items 0.2350 0.4483 0.3000 0.5076

### v2

@@ 5.1. Task Description
Table 8. Task-specific prompts for LLMs as recommender systems, and we categorize them into two kinds of tasks, i.e., accuracy-focused recommendation task and beyond-accuracy recommendation task.
Recommendation Task Task-specific Prompts
Accuracy-focused Recommendation Task
Rating Prediction Based on the user’s historical behavior and item features, predict the rating the user would give to the item (Geng et al., 2022).
Click-through Rate Prediction Given the user’s preference and unpreference, identify whether the user will like the target movie by answering ”Yes.” or ”No.” (Bao et al., 2023b).
Sequential Recommendation A user has interacted with the following items in order, recommend the next item the user will interact with (Hou et al., 2024).
Conversational Recommender System Pretend you are a movie recommender system. I will give you a conversation between a user and you (He et al., 2023).
Beyond-accuracy Recommendation Task
Explainable Recommendation As a movie recommender system, give reasons why the customer watches this movie Return of the Jedi (Luo et al., 2024).
Fairness-aware Recommendation Act as a fair recommender system balancing between popular and less-known movies to ensure fairness (Deldjoo, 2024).
Diversified Recommendation Your task is to re-rank this candidate list and provide a final top-$n$ recommendation list where the goal is to balance relevance and diversity (Carraro and Bridge, 2024).
Task description of recommendation tasks for LLM-RS is a key part of prompts that makes LLM understand the goal of specific recommendation assignments. Typically, based on the different recommendation objectives, existing recommendation tasks can be divided into the following categories: point-wise optimization for a single item (e.g., binary cross entropy loss (Rendle, 2010)), pair-wise optimization for positive and negative items (e.g., Bayesian personalized ranking loss (Rendle et al., 2009)), and list-wise optimization for multiple items (e.g., variational auto-encoders (Liang et al., 2018)). For LLM-RS, we can adapt our work into multiple tasks by changing the description of prompts and the form of candidate items. That is to say, our framework can not only be used for re-ranking (Hou et al., 2024) and retrieving (Li et al., 2023c) tasks, but also support task forms such as point-wise prediction (Kang et al., 2023), pair-wise comparison (Dai et al., 2023a) and list-wise re-ranking formats. As illustrated in Table 8, we list task-specific prompts for LLMs as recommender systems.
$\bullet$ Point-wise recommendation  (Cheng et al., 2016) regards the recommendation task as a binary classification problem, such as the rating scoring task (Geng et al., 2022; Chu et al., 2023) and click-through rate (CTR) predictions (Kang et al., 2023; Bao et al., 2023b). It generates a score or likelihood of preference for a given user-item pair with feature engineering. The recommender system ranks items solely based on the individual characteristics or historical interactions of users without considering the comparison and mutual influence between items in the input ranking list. When utilizing LLMs for point-wise recommendation, the description mainly involves the user-item information with the range of score or answer list to limit output formats.
@@
$\bullet$ List-wise recommendation (Liang et al., 2018) involves optimizing an entire list of recommended items for a user. Instead of considering items individually or in pairs, this method treats the entire recommendation list as a single entity and captures the interior correlations within the list. It aims to create a ranked list of items that collectively maximizes user satisfaction. For LLM-based list-wise recommendations, the prompt contains a list of items and corresponding instructions, eliciting the ability of LLMs to explore the potential relationship within the item sequence (Hou et al., 2024; Ma et al., 2023).
$\bullet$ Matching (Bao et al., 2023a; Li et al., 2023c) in recommendation involves obtaining a small subset of candidates from the entire item pool for users, and candidate items are not provided in prompts at this stage. Due to the lack of domain knowledge in recommender systems, leveraging LLMs for item retrieval requires additional strategies such as model training, item indexing and grounding approaches (Zheng et al., 2024). The prompt for matching tasks includes the description of a user in different forms, e.g., textual profiles and user-item interaction histories, in order to quickly match suitable items.
$\bullet$ Ranking  (Luo et al., 2023) in recommendation involves sorting or ordering items based on the predicted relevance or probability of interest to the target user. It aims to present the most relevant items at the top of the recommendation list. By analyzing historical user-item interactions, user preferences, item features, and the given cadidate item list, it requires LLMs to estimate the relevance of items and determine corresponding positions in the recommendation list (Li et al., 2023j; Luo et al., 2023; Gao et al., 2023).
@@ 5.2. User Interest Modeling
Compared to general tasks solved by LLMs, the characteristic of recommendation tasks lies in the mining and utilization of the user personalized interest (Yao et al., 2023; Shu et al., 2023). Users are influenced by multiple factors when making recommendation decisions, including but not limited to their long-term preferences, short-term intentions, market popular tendencies and occasional environmental biases (Cheng et al., 2016; He et al., 2017). Among them, the most essential aspect is the interest of users, which is the purpose of recommender systems (He et al., 2020; Zhou et al., 2020; Hou et al., 2022). In this section, we first classify the user interest type (Section 5.2.1), outline the user representation forms (Section 5.2.2), and summarize existing modeling methods (Section 5.2.3). Then, we discuss approaches to modeling the user interest (Section 5.2.4), and conduct experiments to provide empirical analysis and key findings.
@@ 5.2.1. User Interest Type
In recommender systems, there are multiple ways to classify the user interest (Rendle et al., 2009; Guo et al., 2017; Zhou et al., 2020). For example, based on the type of feedback between users and the platform, the user interest can be divided into the explicit interest (e.g., ratings) and implicit interest (e.g., clicks). Although the explicit feedback can reflect the true intentions of users, the sparse and expensive data limits its application scenarios. In general, we mainly explore user interest modeling in LLMs under implicit feedback scenarios. Another classification of the user interest is based on the time duration and stability, i.e., short-term interest, long-term interest, and hybrid interest as follows:
$\bullet$ Short-term interest refers to the sudden and accidental intentions or tendencies of personalized users in recent interactions, which is prone to change and can be influenced by environmental factors (Lin et al., 2024a; Hou et al., 2024). That is to say, short-term interest is recent, temporary, and variable (Zhang et al., 2023e).
$\bullet$ Long-term interest refers to the stable preferences of users towards certain content, themes, and elements, which is not easily changed in the short term and will continue to affect recommendation decisions of users (Shu et al., 2023; Wang et al., 2024c). In contrast to short-term intentions, long-term interest is long-term, sustained and stable.
$\bullet$ Hybrid interest means the combination of long-term preferences and short-term intentions (Friedman et al., 2023; Liu et al., 2024d; Li et al., 2023b). On the one hand, short-term intentions are dominated by long-term preferences (Yue et al., 2023). On the other hand, long-term preferences consist of short-term interest across periods (Li et al., 2023i; Qiu et al., 2023).
@@ 5.2.2. User Representation Forms
In LLM-RS, user interest modeling should consider both the input form for the prompt template and the specific storage form for the interest memory (Wang et al., 2024c). Generally, there are three types of input contents for interest modeling in LLMs: historical item lists, interest descriptions and user embeddings. Corresponding to different representation forms of the user interest, there are also various storage forms in the interest memory.
@@
$\bullet$ Historical item lists refer to representing personalized users based on the sequence of historical interacted items, which is widely used in session-based recommendation and sequential recommendation (Hou et al., 2022; Zhou et al., 2020; Kang and McAuley, 2018). For token-based item indexing, the ID sequence of items is used for user modeling in LLMs, similar to traditional sequential recommendation models (Geng et al., 2022). However, without fine-tuning, LLMs cannot recognize the meaning of corresponding IDs. For description-based item indexing, existing researches generally concatenate the attributes (e.g., titles) of items in the temporal order, and then input the item sequence as text descriptions to LLMs as user representations (Li et al., 2023e, b; Hou et al., 2024). Only the item IDs that the user has interacted with need to be stored as the interest memory. The list of items can be stored in the form of numpy or tensor arrays. Despite the simplicity and effectiveness, text attributes such as titles cannot fully represent items with ambiguity (Li et al., 2023e). At the same time, item sequences inevitably contain noise information (Wei et al., 2024b), and limited sequence lengths make it difficult to accurately express the user interest.
@@
$\bullet$ Interest descriptions refer to representing users based on textual descriptions, which is more applicable to the text input of LLMs (Yao et al., 2023). That is to say, we can use natural languages to model the user interest in the form of text and input them into LLMs (Xi et al., 2023). To facilitate the storage and retrieval of textual descriptions, vector stores are commonly used (Touvron et al., 2023b). In LLMs, vector stores in the form of key-value pairs can connect implicit vector embeddings with explicit textual descriptions, improving the retrieval efficiency and comprehension ability (Li et al., 2023c; Petrov and Macdonald, 2023). However, the key lies in how to mine and describe the user interest, and this is what we should make efforts to discuss.
@@
$\bullet$ User embeddings refer to concatenating embeddings of users to the input of LLMs as user representations (Zhang et al., 2023b), which is often used for efficient fine-tuning of language models. In this case, we assign each user a unique ID and corresponding embedding, and add the vector representations of users to the input of LLMs for fine-tuning. At the same time, user embeddings can also be trained from small-scale traditional recommendation models (Zhang et al., 2023c), thereby incorporating collaborative filtering features from other users. However, it is worth noting that vector embeddings of user representations are only suitable for scenarios where the domain knowledge can be injected into LLMs, and the black-box property of embeddings also poses challenges to the explainability.
@@ 5.2.3. Modeling Methods
As illustrated in Fig. 9, we classify methods for modeling interest as i.e., memory-based methods, retrieval-based methods and generation-based methods.
@@
$\bullet$ Memory-based methods assign external memory to users for storing interest-related historical information (Xi et al., 2023; Shu et al., 2023). As for user representations, the encoded interest of users can be obtained from the memory (Zhang et al., 2023c), and LLMs are instructed to utilize the memorized interest with carefully designed prompts. The iterative updating of memories is the key to memory-based methods. As for personalized memory, there are three important operations (Wang et al., 2023e): (1) memory reading is to obtain contents of the interest memory based on the user identifier. (2) Memory writing is to augment new interest of users based on the latest interactions between users and items. (3) Memory reflection is the periodic examination and updating of existing issues in the memory. Strategies for memory reflection include but are not limited to self-summarization, self-correction, and reflection based on user feedback (Wang et al., 2023a). It is memory reflection that makes memory not just a stack of historical records but a summary of user interest (Yao et al., 2023). Note that memory-based methods specifically refer to obtaining contents from the memory without further processing procedures.
@@
$\bullet$ Retrieval-based methods add a module for personalized interest retrieval on top of the memory-based methods, utilizing retrieval strategies to obtain user interest from the personalized memory (Wang et al., 2024c; Lin et al., 2024a; Huang et al., 2023). As for the personalized query, there are generally three ways to form a query for retrieval: (1) candidate items as search queries (Lin et al., 2024a) for relevant items, (2) recently interacted items as search queries for similar items (Zhang et al., 2024a), and (3) the user profile as search queries for personalized items (Wang et al., 2024c). For the criteria of retrieval, there are multiple trade-offs, including dimensions such as relevance, recency, and diversity with respect to the user interest (Wang et al., 2023e).
@@
$\bullet$ Generation-based methods utilize the generation capabilities of LLMs to summarize, infer, and derive comprehensive user interest (Yao et al., 2023; Wang and Lim, 2023; Luo et al., 2023). Generally, a combination of memory-based and retrieval-based methods is required for generation-based methods. For generative language models, the user interest induced by generative retrieval can further stimulate the prompting ability of LLMs (Liu et al., 2023a; Zhang et al., 2023d). The generative approach can also be used for data augmentation, leveraging the general knowledge of LLMs to augment textual descriptions of the user interest. In current recommendations, LLMs mainly utilize the retrieved relevant items to generate summarized interest descriptions, which can also serve as a strategy for the memory reflection (Wang et al., 2024c; Shu et al., 2023).
Figure 9. Comparison of three modeling methods of user interest, i.e., memory-based methods, retrieval-based methods, and generation-based methods. Interest memory consists of global and personalized memory.
[…]
Comparison of three modeling methods of user interest, i.e., (1) memory-based methods, (2) retrieval-based methods, and (3) generation-based methods. Interest memory consists of global and personalized memory.
5.2.4. Research Questions and Observations for Modeling User Interest
@@
In this section, we explore methods for user interest modeling in LLM-RS w.r.t. the following two research questions.
  • RQ1: For retrieval-based modeling methods, how to design query and memory to enrich the expression of user interest?
  • RQ2: How do different methods of user interest modeling affect the recommendation results?
In this section, we conduct experiments for the zero-shot re-ranking task as illustrated in Fig. 2(a), and ChatGPT is employed as the zero-shot re-ranker.
@@
For RQ1, to analyze how to select recent and relevant items as the representation of user interest, we adopt the retrieval-based strategy to retrieve items from the long-term interest. Specifically, following the memory-based and retrieval-based strategy, the personalized query, e.g., candidate items, recently interacted items or the user profile, is encoded into vectors in a unified form. The most relevant items are retrieved from memory based on semantic similarity. Meanwhile, we set weights based on the recency of interactions so that the recent items have a higher probability of being retrieved. Finally, the corresponding retrieved text is obtained based on key-value pairs in the memory. Here, we explore the design of the personalized query and memory of users for item retrieval, and study the influence of different variants as follows.
@@
  • Variants of the personalized query: as for the personalized query, we utilize the interest of users for retrieving items, and compare two strategies for constructing the query.
@@
    • Long- and short-term interest: we first summarize the personalized profile of each user based on all items he or she has interacted, then concatenate the user profile and the summarized text based on 10 recently interacted items for the memory retrieval of the long-term interest. We utilize the long-term user profile and short-term interest for the memory retrieval of the long-term interest.
  • Variants of the memory: as for the personalized memory, we also consider two variants.
    • Global memory: we utilize descriptions of items in the datasets as the memory. Since all users in a dataset share the same item space, the memory w.r.t. item descriptions is not the personalized memory but global memory.
@@
For RQ2, to examine the impact of modeling forms on the user interest, we design 10 modeling forms according to short-term interest, long-term interest and hybrid interest as follows:
  • Variants of short-term interest modeling forms:
    • Recent items: we use titles of the recent 10 items as the short-term interest.
    • Personalized interest of recent items: we construct the personalized memory for each user based on the user-item interactions, and use the personalized descriptions of recently interacted items as the short-term interest.
    • Recent items with personalized interest: we employ both titles of the recent 10 items, and the personalized text of recent items to prompt LLMs as the short-term interest.
    • Recent items with LLM-generated short-term interest: we utilize the least-to-most prompting strategy as mentioned in Section 5, which summarizes the interest text based on the recently interacted 10 items, then concatenate the summarized text and the 10 items.
  • Variants of long-term interest modeling forms:
    • Retrieved items from user history, recent items as query: it retrieves 10 relevant items from the long-term interest. Since recent items have a higher probability to be retrieved.
    • Retrieved personalized interest, recent items as query: it also retrieves 10 relevant items from the long-term history like “Retrieved items from user history, recent items as query”, but this variant uses the personalized descriptions of items from the personalized memory.
    • LLM-generated summarization of retrieved personalized interest: it employs the retrieved personalized text from long-term history in the last variant by summarizing the retrieved 10 texts into one description. Not providing 10 specific items for LLMs, this variant only summarizes the user interest through personalized text.
  • Variants of hybrid interest modeling forms:
    • Recent and retrieved relevant items: it uses both the retrieved 10 items from long-term history and the recent 10 items as the representation of user interest, considering both the long-term and short-term interest. It is worth noting that the retrieved items in this variant do not include the latest 10 items.
    • Retrieved personalized interest, profile and recent items as query: it retrieves recent and relevant items from the long-term interest, and for the query, both the personalized user profile and short-term interest are utilized.
    • Recent items with LLM-generated long-term interest: it concatenates the user profile summarized by LLMs and the recently interacted 10 items for interest modeling.
Table 9. Performance comparison of user interest modeling forms across three modeling methods. The red color indicates short-term interest modeling, blue indicates long-term interest modeling, and purple indicates hybrid interest modeling. Evaluation metrics N@1 and N@10 represent NDCG@1 and NDCG@10, respectively, and the same below. Following the experimental setup outlined in Table 4, we evaluate the zero-shot recommendation performance of ChatGPT on the MovieLens-1M and Amazon Books datasets.
MovieLens-1M Amazon Books
Modeling Form of User Interest N@1 N@10 N@1 N@10
Memory-based Modeling Methods
Recent items 0.1817 0.3985 0.2467 0.4276
Personalized interest of recent items 0.2100 0.4380 0.2500 0.4733
Recent items with personalized interest 0.2350 0.4483 $0$ul0.5076
Retrieval-based Modeling Methods
Retrieved relevant items from user history, recent items as query 0.1950 0.4343 0.2850 0.4650
[…]
Retrieved personalized interest, profile and recent items as query 0.2400 0.4563 0.3300 0.4712
Generation-based Modeling Methods
Recent items with LLM-generated short-term interest 0.2783 0.5099 0.2617 0.4729
Recent items with LLM-generated long-term interest 0.2000 0.4466 0.2500 0.4925
LLM-generated summarization of retrieved personalized interest 0.2000 0.4492 0.2500 0.4872
Table 10. Performance comparison of query and memory designs for retrieval-based methods. We use retrieved relevant items from memory as user interest in prompts. The impact of query forms (short-term interest vs. long- and short-term interest) and memory forms (global memory vs. personalized memory) is analyzed.
Query Memory MovieLens-1M Amazon Books
N@1 N@10 N@1 N@10
@@
[…]
@@
5.2.5. Observations and Discussion